The average time taken to complete an exam, X, follows a normal probability distribution with mean = 60 minutes and standard deviation 30 minutes. What is the probability that a randomly chosen student will take more than 45 minutes to complete the exam?
Answer #1

P(X > 45)

= P((X - )/ > (45 - )/)

= P(Z > (45 - 60)/30)

= P(Z > -0.5)

= 1 - P(Z < -0.5)

= 1 - 0.3085

= 0.6915